An asset was issued 28 months ago. The asset promised just one cash flow of $12000, to be paid to the owner exactly 5 years from the date that the asset was issued. If the required rate of return on this asset is 8%, then what is its present value? Round your answer to the nearest dollar.
Answer #1

Present value FVx(1=(1+r)^n) Here, =2+8/12 A Rate per annum B Number of years C Number of compoundings per per annum A:C Rate

*Please rate thumbs up